IPP Progress Report for the week 2011.12.12 - 2011.12.16

(Up to IPP Progress Reports)

Eugene Magnier

I ran more tests of psphot on MD stacks for Doug to explore the faint end bias. We are finding some curious results, in particular it looks like the current default processing is no longer showing a significant faint end bias. I am running more tests with processing options turned on and off to see what might have caused the effect in the past. The important point, though, is that our current processing seems to have fixed this issue (but it would be nice to know why).

Serge Chastel

Heather Flewelling

  • czar/summitcopy investigator
  • ifaps1 liason
    • answered questions for Klaus and Tom
    • transferred a copy of LAP dvodb to ipp022
  • investigated psf_major changes between tags (yes, there are changes, yes, the current tag is the correct calc)
  • further rsync testing of the big dvodb. It's slow (4-6 hours between minidvodb merges)
  • addstar work continuing (for better ways to handle merges of minidvodbs)

Roy Henderson

  • PSVO:
    • added sub-menus to XML plugin query files
    • wrote recursive function to parse sub-menus
    • plugin files now online so PSVO clients can update dynamically (defaults to local copy if offline)
    • added more queries
    • released new version
  • Loading:
    • usual loading monitoring/czaring
    • processed some of the stragglers from the OLD 3pi processing (some were missing from Gene's backup DVO I had been using)
    • yet more dupes during merge. Worked on a list of batches for Sue to remove.
  • ippToPsps development:
    • datastore delete failing. Spent ages tracking it down. My mistake: fixed.
    • working on clients creating their own scratch Dbs using PID as part of name to avoid conflict and ease of client-start-up
  • looked into providing plots for Heather to asses load on ipp005 during loading

Mark Huber

  • finished stepping through ppStack and working on summary tests.

Bill Sweeney

  • Spent this week implementing the new image processing pipeline stage static sky calibration.
    • new staticskytool modes
    • new script
    • tweaks to recipes and file rules for psastro
    • code to read in Extended source parameters so that they can be saved.
  • Changed the postage stamp server and update processing allow postage stamps to be released without magic censoring. Horray!
  • Changed distribution to allow nightly science data to be released uncesored.
  • Cleaned up distribution mess caused by removing the magic requirement. Old data got queued for distribution when it was no

longer available.

  • Reran M31 raw stage distribution to replace filesets that were damaged by the loss of the ipp064 RAID

Chris Waters

  • Finished testing pattern correction update/background changes. Although the background sigmas and numbers of detected objects do not seem to change much, the jpeg images look improved in the new version.
  • Fixed bug in shuffle code that could allow two instances of a file to exist on one host. Restarted shuffle.
  • Mailed analysis of diskspace results to Gene and Ken.
  • Implemented changes to SweetSpot? nightly science reduction to allow stacks to be made. Have not enabled WSdiffs for this because of a race condition between diff queuing and stack completion.